### v3

Abstract
Connectivity (or equivalently, unweighted maximum flow) is an important measure in graph theory and combinatorial optimization. Given a graph $G$ with vertices $s$ and $t$, the connectivity $\lambda(s,t)$ from $s$ to $t$ is defined to be the maximum number of edge-disjoint paths from $s$ to $t$ in $G$.
Much research has gone into designing fast algorithms for computing connectivities in graphs. Previous work showed that it is possible to compute connectivities for all pairs of vertices in directed graphs with $m$ edges in $\tilde{O}(m^{\omega})$ time [Chueng, Lau, and Leung, FOCS 2011], where $\omega\in[2,2.3716)$ is the exponent of matrix multiplication. For the related problem of computing “small connectivities,” it was recently shown that for any positive integer $k$, we can compute $\min(k,\lambda(s,t))$ for all pairs of vertices $(s,t)$ in a directed graph with $n$ nodes in $\tilde{O}((kn)^{\omega})$ time [Akmal and Jin, ICALP 2023].
In this paper, we present an alternate exposition of these $\tilde{O}(m^{\omega})$ and $\tilde{O}((kn)^{\omega})$ time algorithms, with simpler proofs of correctness. Earlier proofs were somewhat indirect, introducing an elegant but ad hoc “flow vector framework” for showing correctness of these algorithms. In contrast, we observe that these algorithms for computing exact and small connectivity values can be interpreted as testing whether certain generating functions enumerating families of edge-disjoint paths are nonzero. This new perspective yields more transparent proofs, and ties the approach for these problems more closely to the literature surrounding algebraic graph algorithms.
1 Introduction
Many problems in graph algorithms involve quantifying how “connected” different parts of a network are. One concept developed to study these questions is that of connectivity: given a graph $G$ with vertices $s$ and $t$, the connectivity from $s$ to $t$, denoted by $\lambda(s,t)$, is the maximum number of edge-disjoint paths from $s$ to $t$ in $G$. Connectivity is an old and well-studied measure in graph theory, important in computer science because the connectivity $\lambda(s,t)$ is equal to the maximum flow value from $s$ to $t$ in the unweighted graph $G$. Consequently, significant research has gone into designing fast algorithms for computing connectivities.
Suppose the graph $G$ has $n$ vertices and $m$ edges. Computing an individual connectivity value in $G$ is easy: since the maximum flow between a fixed pair of nodes can be found in almost linear time [CKL+23], we know that for any fixed pair of vertices $(s,t)$ in $G$, we can compute $\lambda(s,t)$ in almost-optimal $m^{1+o(1)}$ time. For many applications however, knowing a single connectivity is not so useful, and it is instead far more informative to know the values of multiple connectivities in a graph.
This motivates the All-Pairs Connectivity (APC) problem, where we are tasked with computing connectivities for all pairs of vertices in a given graph. A long line of work recently culminated in a near-optimal $\tilde{O}(n^{2})$ time algorithm for solving APC over undirected graphs [AKL+22]. Throughout this paper, we focus on the general case where $G$ is directed.
All-Pairs Connectivity (APC)
Given a directed graph $G$, compute $\lambda(s,t)$ for all pairs of vertices $(s,t)$ in $G$.
We can of course solve APC naively in $n^{2}m^{1+o(1)}$ time, simply by solving a separate instance of maximum flow for each pair of vertices. In dense graphs, this naive approach is actually the fastest known algorithm for APC! In slightly sparse graphs however, we can do better and solve APC in $\tilde{O}(m^{\omega})$ time [CLL13], where $\omega$ is the exponent of matrix multiplication (i.e., $\omega$ is the smallest positive real such that two $a\times a$ matrices can be multiplied using $a^{\omega+o(1)}$ arithmetic operations). The current fastest algorithms for matrix multiplication imply that $\omega<2.3716$ [WXXZ23]. If $\omega=2$, then the $\tilde{O}(m^{\omega})$ time algorithm is always at least as fast as the naive approach.
The lack of progress in finding faster algorithms for APC has motivated researchers to consider relaxations of APC, including the $k$-Bounded All-Pairs Connectivity ($k$-APC) problem:
$k$-Bounded All-Pairs Connectivity Problem ($k$-APC)
Given a directed graph $G$, compute $\min(k,\lambda(s,t))$ for all pairs of vertices $(s,t)$ in $G$.
The $k$-APC problem is relevant in contexts where knowing the precise connectivity values between “well-connected” nodes is not important, and instead we care more about distinguishing for each pair of vertices whether its connectivity is small or large (where $k$ is our cutoff for what counts as “small” and “large”). Since the connectivity between any pair of nodes in $G$ is at most $n-1$, the $k$-APC problem recovers the general APC problem when $k=n-1$. As $k$ gets smaller, $k$-APC intuitively becomes easier. Indeed, it was recently shown that for any positive integer $k$, $k$-APC can be solved in $\tilde{O}((kn)^{\omega})$ time [AJ23].
In summary, the following results are known for the APC and $k$-APC problems.
Theorem 1.1.
There is an algorithm solving APC in $\tilde{O}(m^{\omega})$ time.
Theorem 1.2.
There is an algorithm solving $k$-APC in $\tilde{O}((kn)^{\omega})$ time.
The algorithms establishing Theorems 1.1 and 1.2 are straightforward. However, the proofs of correctness for these algorithms presented in the literature are not at all obvious, and involve arguments in a somewhat complicated “flow vector framework.”
In this note, we present simpler proofs of correctness for the APC and $k$-APC algorithms from [CLL13] and [AJ23] respectively.
Comparison With Previous Work
The algorithms for APC and $k$-APC from [CLL13] and [AJ23] work in similar ways. Each algorithm first constructs a certain random matrix $M$ whose rows and columns are indexed by edges of $G$. Then, for each pair of vertices $(s,t)$, the algorithms return the value of
$\operatorname{rank}\leavevmode\nobreak\ M[E_{\text{out}}(s),E_{\text{in}}(t)]$
as the answer for that pair, where $E_{\text{out}}(s)$ is the set of edges exiting $s$ and $E_{\text{in}}(t)$ is the set of edges entering $t$. To prove correctness, one simply needs to show that with high probability, for every pair of vertices $(s,t)$ the rank expression above equals $\lambda(s,t)$ or $\min(k,\lambda(s,t))$, depending on whether $M$ was designed to solve APC or $k$-APC respectively.
Previous proofs of correctness for these algorithms use the flow vector framework. In this framework, we fix a source node $s$, and imagine pumping out random vectors along the edges exiting $s$. Intuitively, we let these vectors propagate throughout the graph and use them to assign vectors to each edge of $G$ in a manner that satisfies certain “flow conservation” rules. One can then argue that for any vertex $s$ and edge $e$, the column vector $M[E_{\text{out}}(s),e]$ equals the vector assigned to edge $e$ in $G$. This interpretation of the entries of $M$ then allows one to establish the connection between ranks of submatrices of $M$ and connectivity.
See [AJ23, Section 3] for a more detailed overview of the approach in previous work.
In this work, we present alternate proofs of correctness for these APC and $k$-APC algorithms. Our proofs provide a combinatorial interpretation of determinants of submatrices of $M$ as generating functions enumerating families of edge-disjoint walks in $G$. This approach lets us directly connect $\operatorname{rank}M[E_{\text{out}}(s),E_{\text{in}}(t)]$ to $\lambda(s,t)$, without introducing any of the auxiliary scaffolding of the flow vector framework.
This new perspective yields simpler and more direct proofs of correctness for the APC and $k$-APC algorithms than what was previously presented in the literature. For example, our proof does not even use Menger’s theorem (the fact that $\lambda(s,t)$ is equal to the minimum number of edge deletions needed to disconnect $t$ from $s$), which previous proofs relied on. From a pedagogical perspective, the new proofs are more transparent, making it clear how and why matrix rank relates to connectivities, so that the exposition of the APC and $k$-APC algorithms in this paper should be easier to teach and motivate compared to previous work.
Finally, our generating function approach also yields somewhat more expressive results than the flow vector framework, leading to an easier proof of correctness for the $k$-APC algorithm in particular. Previously in [AJ23, Section 4], to prove correctness of the $k$-APC algorithm the authors had to manually reprove “low-rank” versions of all the flow vector framework results previously shown by [CLL13] for the general APC problem. This is not necessary in our approach: once we establish our generating function for edge-disjoint walks in $G$, some small additional reasoning yields both the APC and $k$-APC algorithms.
The only technical wrinkle in our approach is that our combinatorial view of the problem involves manipulating formal power series. However, as we discuss later in Remark 4.16, even this ingredient can be removed if one only wishes to establish Theorems 1.1 and 1.2 over directed acyclic graphs (which is already an interesting result, perhaps more suitable for teaching these algorithms in a classroom setting).
Organization
In Section 2 we identify notation and assumptions used throughout the paper. In Section 3 we review standard definitions and properties of formal power series. In Section 4, we construct a matrix of formal power series whose entries enumerate families of edge-disjoint walks in a graph. In Sections 5 and 6 we leverage the construction from Section 4 to give simple proofs of Theorems 1.1 and 1.2 respectively. We conclude in Section 7 by mentioning some connections between our arguments and classical results in combinatorics and computer science, and highlighting open problems related to computing connectivities.
2 Preliminaries
General Notation
Given a positive integer $a$, we let $[a]=\left\{1,\dots,a\right\}$ denote the set of the first $a$ consecutive positive integers.
Graph Assumptions and Notation
Throughout, we let $G$ denote the input graph, with $n$ nodes and $m$ edges. We let $V$ and $E$ denote the vertex and edge sets of $G$ respectively. We assume that $G$ is weakly connected (i.e., the underlying undirected graph of $G$ is connected), so that $n-1\leq m$. This is without loss of generality: if $G$ is disconnected, we can solve APC and $k$-APC on $G$ by solving these problems separately on each weakly connected component of $G$. We assume that $G$ is simple (i.e., $G$ does not have self-loops or parallel edges between nodes).
For a directed edge $e=(u,v)\in E$ going from $u$ to $v$, we let $\operatorname{tail}(e)=u$ and $\operatorname{head}(e)=v$ denote the vertices that $e$ exits and enters respectively.
A walk in $G$ is a sequence of edges $W=\langle e_{1},\dots,e_{\ell}\rangle$ such that $\operatorname{head}(e_{j})=\operatorname{tail}(e_{j+1})$ for each $j\in[\ell-1]$. We say $W$ is a walk starting at $e_{1}$ and ending at $e_{\ell}$. We say $W$ is a path if no two of its edges enter the same vertex, and its starting $s=\operatorname{tail}(e_{1})$ and ending $t=\operatorname{head}(e_{\ell})$ vertices are distinct. Such a path from $s$ to $t$ is referred to as an $st$-path.
Finite Field Computation
Throughout, we work over a finite field $\mathbb{F}=\mathbb{F}_{2^{q}}$ of characteristic two. We set $q=\Theta(\log n)$ large enough so that the field has size $2^{q}\geq 12n^{6}$ (looking ahead, this is to ensure that our algorithms work with high probability). We can perform arithmetic operations over this field in $q^{1+o(1)}=\operatorname{poly}(\log n)$ time.
Matrix Notation
Given a matrix $M$, for any row index $i$ and column index $j$ we let $M[i,j]$ be the $(i,j)$ entry of $M$. Given subsets $I$ and $J$ of row and column indices respectively, we let $M[I,J]$ be the submatrix of $M$ restricted to rows in $I$ and columns in $J$. We also let $M[I,\cdot]$ be the submatrix restricted to rows in $I$ and all columns, and $M[\cdot,J]$ be the submatrix on all rows and restricted to columns in $J$. We let $\operatorname{rank}M$ denote the rank of $M$, defined to be largest nonnegative integer $r$ such that $M$ contains an $r\times r$ submatrix with nonzero determinant. When $M$ is a square matrix, we let $\det M$ denote the determinant of $M$, $\operatorname{adj}(M)$ denote the adjugate of $M$, and $M^{-1}$ denote the inverse of $M$ (if it exists).
Matrix Computation
We recall the following well-known results concerning matrix computation.
Proposition 2.1 (Matrix Inversion).
For any positive integer $a$, we can compute the inverse of an $a\times a$ matrix over a field in $O(a^{\omega})$ field operations.
Proposition 2.2 (Matrix Rank).
For any positive integers $a$ and $b$, we can compute the rank of an $a\times b$ matrix over a field in $O(ab^{\omega-1})$ field operations.
Proofs of Propositions 2.1 and 2.2 can be found in [SCWW21] and [IMH82] respectively.
Identity Testing
To prove correctness of the APC and $k$-APC algorithms, we use the fact that random evaluations of a low degree polynomial (or more generally, rational function whose numerator and denominator have low degree) over a large field are nonzero with high probability.
Proposition 2.3.
Let $P$ be a nonzero $r$-variate polynomial of degree at most $d$. Then a uniform random evaluation of $P$ over $\mathbb{F}^{r}$ is nonzero with probability at least $1-d/|\mathbb{F}|$.
For an accessible proof of Proposition 2.3, see [MR95, Theorem 7.2].
Corollary 2.4 (Rational Identity Testing).
Let $R=P/Q$ be a rational function, represented as the ratio of two nonzero polynomials $P$ and $Q$. Suppose $P$ and $Q$ each have degree at most $d$. If we assign each variable of $R$ an independent, uniform random element of $\mathbb{F}$, then $R$ has nonzero evaluation with probability at least $1-2d/|\mathbb{F}|$.
Proof 2.5.
Under random evaluation over $\mathbb{F}$, by Proposition 2.3 and the union bound, $P$ and $Q$ are both nonzero with probability at least $1-2d/|\mathbb{F}|$. So with this probability, the rational function $R=P/Q$ also has nonzero evaluation, as claimed.
3 Power Series Preliminaries
Our algorithms for computing connectivities work by constructing generating functions for families of edge-disjoint walks. These generating functions involve infinite sums, so in this section we review properties of formal power series, a generalization of polynomials which allow for infinite sums. The results we review are simple, and mostly involve observing that basic facts which hold for polynomials still hold in the infinite case of formal series.
We also note that if one is interested in solving APC and $k$-APC only in the special case of directed acyclic graphs, then it suffice to work with polynomials (no formal power series are needed) and this section can be skipped. We discuss this simplification for acyclic graphs in detail later on in Remark 4.16.
Fix a finite set $J$, and consider the set of variables $\left\{x_{j}\right\}_{j\in J}$ indexed by $J$. A polynomial is a finite linear combination of products of these variables. A formal power series is simply a generalization of polynomials which allows for infinite sums.
Let $\mathbb{N}^{J}$ be the set of all sequences of nonnegative integers indexed by $J$. Given $\boldsymbol{d}\in\mathbb{N}^{J}$, we let $d_{j}$ denote the $j^{\text{th}}$ element of $\boldsymbol{d}$ for each $j\in J$. Then a formal power series $F$ is identified by a sequence of coefficients $a_{\boldsymbol{d}}$ in $\mathbb{F}$, one for each $\boldsymbol{d}\in\mathbb{N}^{J}$, and we write
$F=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}a_{\boldsymbol{d}}\prod_{j\in J}x_{j}^{d_{j}}.$
We let $\boldsymbol{0}$ denote the all-zeros sequence in $\mathbb{N}^{J}$, and say $a_{\boldsymbol{0}}$ is the constant term of $F$. In general, given $\boldsymbol{d}\in\mathbb{N}^{J}$, the monomial corresponding to $\boldsymbol{d}$ in $F$ (if it appears with nonzero coefficient $a_{\boldsymbol{d}}\neq 0$) is said to have degree
$\sum_{j\in J}d_{j}.$
Given formal series
$F=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}a_{\boldsymbol{d}}\prod_{j\in J}x_{j}^{d_{j}}\quad\text{and}\quad H=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}b_{\boldsymbol{d}}\prod_{j\in J}x_{j}^{d_{j}}$
we define their sum
$F+H=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}\left(a_{\boldsymbol{d}}+b_{\boldsymbol{d}}\right)\prod_{j\in J}x_{j}^{d_{j}}$
and product
$F\cdot H=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}\left(\sum_{\begin{subarray}{c}\boldsymbol{d_{1}},\boldsymbol{d_{2}}\in\mathbb{N}^{J}\\ \boldsymbol{d_{1}}+\boldsymbol{d_{2}}=\boldsymbol{d}\end{subarray}}a_{\boldsymbol{d_{1}}}b_{\boldsymbol{d_{2}}}\right)\prod_{j\in J}x_{j}^{d_{j}}$ (1)
in the natural way, generalizing arithmetic over polynomials. These operations make the set of polynomials over $\mathbb{F}$ a subring of the ring of formal power series (where the additive and multiplicative identities are the constant polynomials 00 and $1$ respectively).
Proposition 3.1 (Power Series Inversion).
If $F$ is a formal power series with constant term $1$, then there is a unique formal series $H=F^{-1}$ such that $F\cdot H=1$. Moreover, the constant term of $H$ equals $1$.
Proof 3.2.
Suppose
$F=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}a_{\boldsymbol{d}}\prod_{j\in J}x_{j}^{d_{j}}.$
We define the sequence $b_{\boldsymbol{d}}$ of coefficients in $\mathbb{F}$ for all $\boldsymbol{d}\in\mathbb{N}^{J}$ inductively, by setting $b_{\boldsymbol{0}}=1$, and taking
$b_{\boldsymbol{d}}=-\left(\sum_{\boldsymbol{d^{\prime}}\prec\,\boldsymbol{d}}a_{\boldsymbol{d}-\boldsymbol{d^{\prime}}}b_{\boldsymbol{d^{\prime}}}\right)$ (2)
where $\boldsymbol{d}^{\prime}\prec\boldsymbol{d}$ means that $\boldsymbol{d^{\prime}}\in\mathbb{N}^{J}$ is componentwise less than or equal to $\boldsymbol{d}$, and $\boldsymbol{d^{\prime}}\neq\boldsymbol{d}$.
Then if we set
$H=\sum_{\boldsymbol{d}\in\mathbb{N}^{J}}b_{\boldsymbol{d}}\prod_{j\in J}x_{j}^{d_{j}}$
it follows from the definition of multiplication in Equation 1, the relationship from Equation 2, and the fact that $a_{\boldsymbol{0}}=b_{\boldsymbol{0}}=1$, that we have
$F\cdot H=1.$
This inverse $H$ is unique, because if another formal series $H^{\prime}$ satisfies $F\cdot H^{\prime}=1$, then the constant term of $H^{\prime}$ is $1$ since the product of the constant terms of $F$ and $H^{\prime}$ are $1$, and
$H^{\prime}=H^{\prime}\cdot 1=H^{\prime}\cdot(F\cdot H)=(H^{\prime}\cdot F)\cdot H=1\cdot H=H.$
Thus $F$ has a unique multiplicative inverse.
Matrices of Formal Series
Throughout this paper, we consider matrices whose entries are formal power series. Such matrices naturally arise when computing the inverses of polynomial matrices.
Proposition 3.3 (Geometric Series Formula).
Suppose $X$ is a square matrix with polynomial entries such that every entry of $X$ has constant term equal to zero. Then
$(I-X)^{-1}=\sum_{\ell=0}^{\infty}X^{\ell}.$ (3)
Proof 3.4.
Since every entry of $X$ has zero constant term, every nonzero entry of $X^{\ell}$ has degree at least $\ell$. Consequently, the infinite sum from the right-hand side of Equation 3 is well-defined, because for any $\boldsymbol{d}\in\mathbb{N}^{J}$, only finitely many terms contribute to the coefficient of
$\prod_{j\in J}x_{j}^{d_{j}}$
in each entry of the sum. It suffices to prove that the product
$(I-X)\left(\sum_{\ell=0}^{\infty}X^{\ell}\right)$ (4)
is equal to the identity matrix.
For any fixed integer $d\geq 0$, let $M_{d}$ be the matrix from Equation 4 with entries restricted to terms of degree at most $d$. Then since nonzero entries of $X^{\ell}$ have degree at least $\ell$, we see that $M_{d}$ is equal to the matrix
$(I-X)\left(\sum_{\ell=0}^{d}X^{\ell}\right)=\sum_{\ell=0}^{d}\left(X^{\ell}-X^{\ell+1}\right)=I-X^{d+1}$
with entries restricted to terms of degree at most $d$. Since every nonzero entry of $X^{d+1}$ has degree greater than $d$, we see that $M_{d}=I$ is the identity matrix. Since this equation holds for every fixed $d\geq 0$, the product in Equation 4 equals the identity matrix as claimed.
4 Enumeration with Edge-Adjacency Matrices
In this section we construct a matrix $\Gamma$ of formal power series, whose rows and columns are indexed by edges of $G$. The matrix $\Gamma$ will be designed to have the special property that for any equal-size subsets of edges $S,T\subseteq E$, the determinant
$\det\leavevmode\nobreak\ \Gamma[S,T]$
is nonzero as a formal series if and only if there are edge-disjoint paths connecting the edges in $S$ to the edges in $T$. Since connectivity is defined in terms of edge-disjoint paths, intuitively our construction of $\Gamma$ will let us solve the APC and $k$-APC problems in Sections 5 and 6 by performing certain matrix computations.
For every pair of edges $(e,f)$ in $G$ such that $\operatorname{head}(e)=\operatorname{tail}(f)$ (i.e., edge $e$ enters the vertex that edge $f$ exits), we introduce an indeterminate variable $x_{ef}$.
Let $X$ be the $m\times m$ matrix with rows and columns indexed by edges of $G$, such that for each pair of edges $(e,f)$ in $G$, we have
$X[e,f]=\begin{cases}x_{ef}&\text{if }\operatorname{head}(e)=\operatorname{tail}(f)\\ 0&\text{otherwise.}\end{cases}$ (5)
We enumerate walks not by counting their number, but by assigning each walk a monomial weight, that records information about the edges traversed in the walk. Enumeration for our purposes corresponds to summing the weights of all walks (or collections of walks) in a certain family of interest.
Given a walk $W=\langle e_{1},\dots,e_{\ell}\rangle$, viewed as a sequence of edges $e_{j}$, we let the weight
$\xi(W)=\prod_{j=1}^{\ell-1}x_{e_{j}e_{j+1}}$
of $W$ be the monomial $\xi(W)$ recording all pairs of consecutive edges traversed by $W$. By convention, we assign a walk $W$ of length one (i.e., a single edge) weight $\xi(W)=1$. More generally, given a collection of walks $\mathcal{C}=\langle W_{1},\dots,W_{r}\rangle$ we let the weight
$\xi(\mathcal{C})=\prod_{j=1}^{r}\xi(W_{j})$
of $\mathcal{C}$ be the product of the weights of the individual walks.
Enumerating Walks
Given edges $e,f\in E$ and an integer $\ell\geq 1$, let $\mathcal{W}_{\ell}(e,f)$ denote the set of all walks beginning at $e$ and ending at $f$ of length $\ell$. One way of interpreting the definition of $X$ from Equation 5 is that the $(e,f)$ entry of $X$ enumerates all walks of length two from $e$ to $f$ in $G$. These are precisely the walks in $\mathcal{W}_{2}(e,f)$. The next result observes that higher powers of $X$ enumerate walks of longer lengths in $G$.
Proposition 4.1.
For any edges $e,f\in E$ and integer $\ell\geq 0$, we have
$X^{\ell}[e,f]=\sum_{W\in\mathcal{W}_{\ell+1}(e,f)}\xi(W).$
Proof 4.2.
By expanding out the definition of matrix multiplication, we see that
$X^{\ell}[e,f]=\sum_{\begin{subarray}{c}e_{0},\dots,e_{\ell}\in E\\ e_{0}=e\\ e_{\ell}=f\end{subarray}}\prod_{j=0}^{\ell-1}X[e_{j},e_{j+1}].$
By definition, $X[e_{j},e_{j+1}]=x_{e_{j}e_{j+1}}$ if we can step from $e_{j}$ to $e_{j+1}$ in $G$, and is zero otherwise. Thus, the product
$\prod_{j=0}^{\ell-1}X[e_{j},e_{j+1}]$
is nonzero if and only if $W=\langle e_{0},\dots,e_{\ell}\rangle$ is a walk of length $(\ell+1)$ in $G$. In this case,
$\prod_{j=0}^{\ell-1}X[e_{j},e_{j+1}]=\prod_{j=0}^{\ell-1}x_{e_{j}e_{j+1}}=\xi(W)$
so we have
$X^{\ell}[e,f]=\sum_{W\in\mathcal{W}_{\ell+1}(e,f)}\xi(W)$
as claimed.
Corollary 4.3 (Enumerating Walks).
For any edges $e,f\in E$, we have
$(I-X)^{-1}[e,f]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,f)}\xi(W)\right).$
Proof 4.4.
This result follows by combining the geometric series formula from Proposition 3.3 with the enumerative property of powers of $X$ from Proposition 4.1.
Enumerating Edge-Disjoint Walks
Given subsets of edges $S,T\subseteq E$ of equal size $|S|=|T|=r\geq 1$ and an integer $\ell\geq 1$, we define $\mathcal{F}_{\ell}(S,T)$ to be the family of collections of $r$ walks of total length $\ell$, beginning at different edges of $S$ and ending at different edges of $T$. If we fix some ordering $e_{1},\dots,e_{r}$ of the edges in $S$, then we can view each element of $\mathcal{F}_{\ell}(S,T)$ as a sequence of walks $\langle W_{1},\dots,W_{r}\rangle$ satisfying the properties that each $W_{i}$ begins at $e_{i}$ and ends at some edge of $T$, the $W_{i}$ walks all end at distinct edges of $T$, and the sum of the lengths of the $W_{i}$ walks is $\ell$.
Furthermore, let $\mathcal{D}_{\ell}(S,T)\subseteq\mathcal{F}_{\ell}(S,T)$ be the family of collections of $r$ edge-disjoint walks from $S$ to $T$ of total length $\ell$.
Corollary 4.3 shows that entries of $\Gamma=(I-X)^{-1}$ enumerate walks in $G$. The following result uses this fact to show that determinants of submatrices of $\Gamma$ enumerate collections of walks in $G$, beginning and ending at different edges. This is a simple observation, following immediately from the definition of the determinant. The proof appears somewhat long only because we spell out the details of each step in the calculation.
Lemma 4.5 (Arbitrary Walks).
For any equal-size subsets of edges $S,T\subseteq E$, we have
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
Proof 4.6.
For convenience, write $\Gamma=(I-X)^{-1}$.
Let $\mathfrak{S}(S,T)$ be the set of all bijections from $S$ to $T$.
By the definition of the determinant, we have
$\det\Gamma[S,T]=\sum_{\pi\in\mathfrak{S}(S,T)}\prod_{e\in S}\Gamma[e,\pi(e)].$ (6)
Note that we do not include a factor for the sign of $\pi$ in the above equation, because we are working over a field of characteristic two.
By Corollary 4.3, for each $e\in S$ we have
$\Gamma[e,\pi(e)]=\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e,\pi(e))}\xi(W)\right).$ (7)
Write $S=\left\{e_{1},\dots,e_{r}\right\}$, where $r=|S|=|T|$.
By multiplying the above equation over all choices of $e\in S$, we have
$\prod_{e\in S}\Gamma[e,\pi(e)]=\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\xi(W)\right)\right).$ (8)
Now, let $\mathcal{L}$ be the set of all $r$-tuples $(\ell_{1},\dots,\ell_{r})$ of positive integers summing to
$\ell_{1}+\dots+\ell_{r}=\ell.$
If we expand out the product on the right-hand side of Equation 8 and group terms according to the total length of the walks they come from, we obtain
$\prod_{i=1}^{r}\left(\sum_{\ell=0}^{\infty}\left(\sum_{W\in\mathcal{W}_{\ell+1}(e_{i},\pi(e_{i}))}\xi(W)\right)\right)=\sum_{\ell=1}^{\infty}\left(\sum_{\begin{subarray}{c}(\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$
To clarify the expression above: in the right-hand side of the above equation, the first inner summation is over all choices of positive integers $\ell_{1},\dots,\ell_{r}$ which sum to $\ell$, and choices of walks $W_{1},\dots,W_{r}$ where $W_{i}$ is a walk of length $\ell_{i}$ from $e_{i}$ to $\pi(e_{i})$. This is simply the result of distributing the product over $i\in[r]$ on the left-hand side of the equation over the sum of walks of all possible lengths from $e_{i}$ to $\pi(e_{i})$.
By chaining the above equation together with Equations 6, 7 and 8, and interchanging summation, we get that
$\det\leavevmode\nobreak\ \Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\pi\in\mathfrak{S}(S,T)}\sum_{\begin{subarray}{c}\\ (\ell_{1},\dots,\ell_{r})\in\mathcal{L}\\ W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))\end{subarray}}\prod_{i=1}^{r}\xi(W_{i})\right).$ (9)
To simplify Equation 9, observe that for any choice of bijection $\pi\in\mathfrak{S}(S,T)$, lengths $(\ell_{1},\dots,\ell_{r})\in\mathcal{L}$, and walks $W_{i}\in\mathcal{W}_{\ell_{i}}(e_{i},\pi(e_{i}))$, the collection $\langle W_{1},\dots,W_{r}\rangle$ is a sequence of walks from $S$ to $T$ of total length $\ell$. Conversely, any collection $\mathcal{C}\in\mathcal{F}_{\ell}(S,T)$ has walks whose lengths sum up to $\ell$, and corresponds to a unique bijection $\pi\in\mathfrak{S}(S,T)$, obtained by checking which starting edges in $S$ are connected to which ending edges in $T$ by walks in $\mathcal{C}$.
Thus, the inner nested summation above is equivalent to a single sum over all collections of walks in $\mathcal{F}_{\ell}(S,T)$. Since the weight of a collection $\mathcal{C}=\langle W_{1},\dots,W_{r}\rangle$ is precisely
$\xi(\mathcal{C})=\prod_{i=1}^{r}\xi(W_{i}),$
the discussion from the previous paragraph together with Equation 9 implies that
$\det\Gamma[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})\right)$
which proves the desired result.
We now observe that the determinant sieves out collections of intersecting walks, so that only edge-disjoint families of walks are included in its enumeration.
Lemma 4.7 (Intersecting Walks Cancel).
For any equal-size subsets of edges $S,T\subseteq E$ and integer $\ell\geq 1$, we have
$\sum_{\mathcal{C}\in\mathcal{F}_{\ell}(S,T)}\xi(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C}).$
Proof 4.8.
Fix $S,T\subseteq E$ and integer $\ell\geq 1$. Let $r=|S|=|T|$.
For convenience, abbreviate $\mathcal{F}=\mathcal{F}_{\ell}(S,T)$ and $\mathcal{D}=\mathcal{D}_{\ell}(S,T)$. Let $\mathcal{S}=\mathcal{F}\setminus\mathcal{D}$ be the family of all collections of $r$ walks beginning at different edges of $S$ and ending at different edges of $T$, such that at least two walks in the collection intersect at an edge. By definition we have
$\sum_{\mathcal{C}\in\mathcal{F}}\xi(\mathcal{C})=\sum_{\mathcal{C}\in\mathcal{D}}\xi(\mathcal{C})+\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C}).$
So to prove the claim, it suffices to show that
$\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C})$
is the zero polynomial. We prove this by pairing up collections $\mathcal{C}$ in $\mathcal{S}$ of equal weight $\xi(\mathcal{C})$, and observing that contributions from such collections vanish modulo two.
$e_{1}$$e_{2}$${\color[rgb]{0.36,0.22,0.33}\definecolor[named]{pgfstrokecolor}{rgb}{0.36,0.22,0.33}g}$$e_{3}$$e_{4}$$f_{1}$$f_{2}$$f_{3}$$f_{4}$$e_{1}$$e_{2}$${\color[rgb]{0.36,0.22,0.33}\definecolor[named]{pgfstrokecolor}{rgb}{0.36,0.22,0.33}g}$$f_{3}$$f_{4}$$f_{1}$$f_{2}$$e_{3}$$e_{4}$
Figure 1: Given walks $W_{i}=\langle e_{1},e_{2},g,e_{3},e_{4}\rangle$ and $W_{j}=\langle f_{1},f_{2},g,f_{3},f_{4}\rangle$ overlapping at $g$, we can swap their suffixes to produce walks $W^{\prime}_{i}=\langle e_{1},e_{2},g,f_{3},f_{4}\rangle$ and $W^{\prime}_{j}=\langle f_{1},f_{2},g,e_{3},e_{4}\rangle$ which still overlap at $g$. The weights for both pairs $\xi(W_{i},W_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})$ and $\xi(W^{\prime}_{i},W^{\prime}_{j})=(x_{e_{1}e_{2}}x_{e_{2}g}x_{gf_{3}}x_{f_{3}f_{4}})\cdot(x_{f_{1}f_{2}}x_{f_{2}g}x_{ge_{3}}x_{e_{3}e_{4}})$ agree because each pair traverses the same multiset of consecutive pairs of edges.
Fix an ordering $e_{1},\dots,e_{r}$ of the edges in $S$. Take any $\mathcal{C}=\langle W_{1},\dots,W_{r}\rangle\in\mathcal{S}$, with the walks ordered so that $W_{i}$ begins at edge $e_{i}$. By assumption, at least two walks in $\mathcal{C}$ overlap at an edge. Let $i\in[r]$ be the smallest index such that $W_{i}$ intersects some other walk in $\mathcal{C}$ at an edge. Let $e$ be the first edge in $W_{i}$ which is contained in another walk of $\mathcal{C}$. Let $j\in[r]$ be the smallest index $j>i$ such that $W_{j}$ overlaps with $W_{i}$ at edge $e$.
We can split the walk $W_{i}$ uniquely
$W_{i}=A_{i}\diamond B_{i}$
as the concatenation of a prefix walk $A_{i}$ not including edge $e$, and a suffix walk $B_{i}$ beginning with edge $e$. We can similarly split $W_{j}$ uniquely
$W_{j}=A_{j}\diamond B_{j}$
as the concatenation of a prefix $A_{j}$ not including $e$, and a suffix $B_{j}$ beginning with $e$.
Now, define walks
$W^{\prime}_{i}=A_{i}\diamond B_{j}\quad\text{and}\quad W^{\prime}_{j}=A_{j}\diamond B_{i}$
by swapping the suffixes of $W_{i}$ and $W_{j}$. An example of this operation is depicted in Figure 1. For all $l\in[r]$ with $l\not\in\left\{i,j\right\}$, set $W^{\prime}_{l}=W_{l}$. Define a new collection of walks
$\mathcal{C}^{\prime}=\langle W_{1}^{\prime},\dots,W^{\prime}_{r}\rangle$
by replacing $W_{i}$ and $W_{j}$ in $\mathcal{C}$ with $W^{\prime}_{i}$ and $W^{\prime}_{j}$ respectively.
Since $W_{i}$ and $W_{j}$ end at different edges of $T$, we know that $W^{\prime}_{i}\neq W_{i}$ and $W^{\prime}_{j}\neq W_{j}$. This shows that $\mathcal{C}^{\prime}\neq\mathcal{C}$. Since the walks in $\mathcal{C}^{\prime}$ still begin at different edges of $S$ and end at different edges of $T$, $\mathcal{C}^{\prime}\in\mathcal{F}$. Moreover, since $W^{\prime}_{i},W^{\prime}_{j}$ overlap at an edge, we have $\mathcal{C}^{\prime}\not\in\mathcal{D}$.
Thus $\mathcal{C}^{\prime}\in\mathcal{S}$.
Additionally, we claim that if we apply the above suffix swapping procedure (which we used to go from $\mathcal{C}$ to $\mathcal{C}^{\prime}$) to the collection $\mathcal{C}^{\prime}$, we recover $\mathcal{C}$.
Indeed, for all $l\in[r]$ with $l<i$, the walk $W^{\prime}_{l}=W_{l}$ does not intersect any other walk in $\mathcal{C}$ at an edge, by the definition of $i$. Since the multiset of edges traversed by walks in $\mathcal{C}\setminus\left\{W_{l}\right\}$ and $\mathcal{C}\setminus\left\{W^{\prime}_{l}\right\}$ are the same, this means that $W^{\prime}_{l}$ does not intersect any other walk in $\mathcal{C}^{\prime}$ at an edge either. So $i$ is the smallest index in $[r]$ such that $W^{\prime}_{i}$ intersects some other walk in $\mathcal{C}^{\prime}$ at an edge. Since the prefixes of $W^{\prime}_{i}$ and $W_{i}$ before edge $e$ are the same, we see that $e$ is also the first edge in $W^{\prime}_{i}$ which is contained in another walk of $\mathcal{C}^{\prime}$. Then because $W_{j}^{\prime}$ traverses edge $e$, and $W_{l}^{\prime}=W_{l}$ for all $l\not\in\left\{i,j\right\}$, we get that $j>i$ is the smallest index such that $W_{j}^{\prime}$ overlaps with $W^{\prime}_{i}$ at edge $e$. Then, when we swap the suffixes of $W^{\prime}_{i}$ and $W^{\prime}_{j}$ after the first appearance of $e$ on these walks, we recover $W_{i}$ and $W_{j}$ respectively, and so applying the suffix swapping procedure to $\mathcal{C}^{\prime}$ produces the original collection $\mathcal{C}$ as claimed.
So, the suffix swapping routine described above partitions $\mathcal{S}$ into distinct pairs.
Suppose $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are paired up by the suffix swapping argument. Then $\mathcal{C}$ and $\mathcal{C}^{\prime}$ traverse the same multiset of consecutive pairs of edges. Thus these collections
$\xi(\mathcal{C})=\xi(\mathcal{C}^{\prime})$
have the same weight. Since we work over a field of characteristic two, the above equation implies that each pair $(\mathcal{C},\mathcal{C}^{\prime})$ of collections mapped to each other by suffix swapping satisfies
$\xi(\mathcal{C})+\xi(\mathcal{C}^{\prime})=0.$
Since $\mathcal{S}$ is partitioned into such pairs, we have
$\sum_{\mathcal{C}\in\mathcal{S}}\xi(\mathcal{C})=0.$
Together with the discussion from the beginning of the proof, this proves the claim.
Remark 4.9 (Characteristic Two is Unnecessary).
In the proof of Lemma 4.7, our argument used the fact that we work over a field of characteristic two. This restriction on the characteristic is only included for the sake of simplicity, and is not necessary for Lemma 4.7 to hold. Indeed, if we instead worked over a field of odd characteristic, then all that changes is that terms in the expansion of the determinant come with a sign, and we can now pair up and cancel terms with opposite signs.
Corollary 4.10 (Edge-Disjoint Walks).
For any equal-size subsets of edges $S,T\subseteq E$, we have
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$
Proof 4.11.
This follows by combining Lemmas 4.5 and 4.7.
Lemma 4.12 (Edge-Disjoint Walks $\Rightarrow$ Edge-Disjoint Paths).
Let $S,T\subseteq E$ be subsets of edges of size $|S|=|T|=r$. If the graph $G$ contains $r$ edge-disjoint walks from $S$ to $T$, then $G$ also contains $r$ edge-disjoint paths from $S$ to $T$
Proof 4.13.
Let $\langle W_{1},\dots,W_{r}\rangle$ be a collection of edge-disjoint walks from $S$ to $T$ in $G$. For each index $i\in[r]$, let $e_{i}$ and $f_{i}$ be the first and last edges of $W_{i}$ respectively. Note that under these definitions, we have $S=\left\{e_{1},\dots,e_{r}\right\}$ and $T=\left\{f_{1},\dots,f_{r}\right\}$.
For each $i\in[r]$, let $G_{i}$ be the subgraph of $G$ including only the edges traversed by $W_{i}$. Let $P_{i}$ be a shortest path from $e_{i}$ to $f_{i}$ in $G_{i}$. These paths are edge-disjoint, since they live in subgraphs on disjoint sets of edges. Thus $\langle P_{1},\dots,P_{r}\rangle$ is a collection of $r$ edge-disjoint paths from $S$ to $T$ in $G$, as desired.
Corollary 4.14.
Let $S,T\subseteq E$ be subsets of edges of size $|S|=|T|=r$. Then
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]$
is a nonzero formal power series if and only if $G$ contains $r$ edge-disjoint paths from $S$ to $T$.
Proof 4.15.
Suppose $\mathcal{C}=\langle P_{1},\dots,P_{r}\rangle$ is a collection of $r$ edge-disjoint paths from $S$ to $T$ in $G$. Then the term $\xi(\mathcal{C})$ occurs in the expansion of
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]$ (10)
given by Corollary 4.10. Moreover, any collection of paths $\mathcal{C}^{\prime}\neq\mathcal{C}$ from $S$ to $T$ has weight $\xi(\mathcal{C}^{\prime})\neq\xi(\mathcal{C})$, because $\mathcal{C}$ consists of edge-disjoint paths (so looking at the variables appearing in $\xi(\mathcal{C})$, we can recover $\mathcal{C}$ uniquely). Hence, no other term from the summation in Corollary 4.10 produces the same monomial $\xi(\mathcal{C})$. So $\xi(\mathcal{C})$ appears in Equation 10 with nonzero coefficient, which implies that the determinant from Equation 10 is a nonzero formal power series.
Suppose now that $G$ does not contain $r$ edge-disjoint paths from $S$ to $T$. The contrapositive of Lemma 4.12 implies that $G$ does not contain $r$ edge-disjoint walks from $S$ to $T$ either. Then Corollary 4.10 implies that Equation 10 is the zero polynomial. This proves the claim.
Remark 4.16 (Directed Acyclic Graphs).
The arguments in this section work with formal power series because when designing a generating function for edge-disjoint walks, we naturally run into infinite sums. From a pedagogical perspective, dealing with these infinite objects may make teaching these algorithms appear somewhat difficult (say, in an undergraduate course). One way to avoid this issue is to focus on the APC and $k$-APC problems in the special case of directed acyclic graphs (DAGs).
DAGs are an interesting case for APC and $k$-APC, since we do not know of algorithms solving these problems on DAGs faster than the case of general directed graphs, and the best conditional lower bounds we have for these problems hold in the case of DAGs [AGI+19].
Focusing on the setting of DAGs leads to two simplifications.
First, any walk in a DAG has length at most $n-1$. Hence over DAGs, $X^{\ell}$ is the all-zeros matrix for $\ell\geq n$, so Proposition 3.3 shows that $\Gamma=(I-X)^{-1}$ is a matrix whose entries are polynomials. Thus, determinants of submatrices of $\Gamma$ are just polynomials, instead of formal power series involving infinite sums. Second, any walk in a DAG is a path. Thus we can skip the step in Lemma 4.12 of going from edge-disjoint walks to edge-disjoint paths.
5 Connectivity
5.1 Random Evaluation
Let $X$ be the symbolic edge-adjacency matrix defined in Section 4.
For all pairs of edges $(e,f)$ in $G$, we introduce independent, uniform random values $a_{ef}$ over $\mathbb{F}$. Let $A$ be the matrix obtained from $X$ by evaluating each variable $x_{ef}$ at $a_{ef}$. That is, $A$ is the random $m\times m$ edge-adjacency matrix of $G$, defined by taking
$A=\begin{cases}a_{ef}&\text{if }\operatorname{head}(e)=\operatorname{tail}(f)\\ 0&\text{otherwise.}\end{cases}$
Lemma 5.1.
Let $S,T\subseteq E$ be subsets of edges of size $|S|=|T|=r\leq n-1$. Then
$\det\leavevmode\nobreak\ (I-A)^{-1}[S,T]$
is nonzero with probability at least $1-1/n^{3}$ if and only if $G$ contains $r$ edge-disjoint paths from $S$ to $T$.
Proof 5.2.
By Corollary 4.14, the determinant
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]$ (11)
is a nonzero formal power series if and only if $G$ contains $r$ edge-disjoint paths from $S$ to $T$.
So suppose $G$ does not contain $r$ edge-disjoint paths from $S$ to $T$. Then Equation 11 is the zero polynomial, so its random evaluation
$\det\leavevmode\nobreak\ (I-A)^{-1}[S,T]$
must vanish as claimed.
Otherwise, suppose $G$ does contain $r$ edge-disjoint paths from $S$ to $T$. Then Equation 11 is a nonzero formal power series.
By the formula for the inverse of a matrix, we know that
$(I-X)^{-1}[S,T]=\frac{(\operatorname{adj}(I-X))[S,T]}{\det\leavevmode\nobreak\ (I-X)}.$ (12)
Since $(I-X)$ has ones along the diagonal and its other entries have constant term zero, we know that $\det(I-X)$ is a polynomial with constant term $1$, so by Proposition 3.1 the multiplicative inverse of $\det(I-X)$ is a well-defined formal power series. Thus, Equation 12 can be viewed as an equality between two matrices of formal power series.
For convenience, write $Q=\det(I-X)$. Since $S$ and $T$ are sets of size $r$, by linearity of the determinant we have
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\frac{\det\leavevmode\nobreak\ (\operatorname{adj}(I-X))[S,T]}{Q^{r}}.$ (13)
By assumption, the left-hand side of Equation 13 is nonzero. Consequently, the numerator
$\det\leavevmode\nobreak\ (\operatorname{adj}(I-X))[S,T]$
on the right-hand side of Equation 13 must be a nonzero polynomial. Moreover, since each entry of $X$ has degree at most $1$, each entry of $\operatorname{adj}(I-X)$ has degree less than $m$, so this numerator polynomial has overall degree less than $rm$. Similarly, in the previous discussion we observed that $Q$ is a polynomial with constant term $1$, so the denominator
$Q^{r}=\left(\det\leavevmode\nobreak\ (I-X)\right)^{r}$
has constant term $1$ and is thus a nonzero polynomial as well. Since each entry of $X$ has degree at most $1$, this denominator polynomial has degree at most $rm$.
The previous paragraph shows that the expression from Equation 13 is the ratio of two nonzero polynomials, each with degree at most $rm<nm$.
Then by rational identity testing (Corollary 2.4), the random evaluation
$\det\leavevmode\nobreak\ (I-A)^{-1}[S,T]$
of Equation 11 over $\mathbb{F}$ is nonzero with probability at least $1-2nm/2^{q}$.
Since we picked $q$ such that $2^{q}\geq 12n^{6}>n^{3}\cdot(2nm)$, the desired result follows.
Lemma 5.3 (Connectivity via Rank).
With high probability, for all $s,t\in V$ we have
$\lambda(s,t)=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)].$
Proof 5.4.
Fix a pair of vertices $(s,t)$. Abbreviate $\lambda=\lambda(s,t)$.
By Lemma 5.1 and the definition of connectivity, with probability at least $1-1/n^{3}$, $\lambda$ is the largest nonnegative integer such that there exist subsets $S\subseteq E_{\text{out}}(s)$ and $T\subseteq E_{\text{in}}(t)$ of size $\lambda$ with the property that
$\det\leavevmode\nobreak\ (I-A)^{-1}[S,T]$ (14)
is nonzero. By definition, this is the rank of $(I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$, so
$\lambda=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$
with probability at least $1-1/n^{3}$ for our fixed pair of vertices $(s,t)$.
Then by the union bound, with probability at least $1-1/n$ we have
$\lambda(s,t)=\operatorname{rank}\leavevmode\nobreak\ (I-A)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$
for all pairs of vertices $(s,t)$ in $G$, as desired.
5.2 The Algorithm
See 1.1
Proof 5.5.
By Lemma 5.3, Algorithm 1 solves APC correctly with high probability. It remains to bound the runtime of the algorithm.
In step 1 of Algorithm 1, we compute $M$ by inverting an $m\times m$ matrix. By Proposition 2.1, this matrix inversion can be performed in $\tilde{O}(m^{\omega})$ time.
In step 2 of Algorithm 1, we compute $\lambda(s,t)$ for each pair of vertices $(s,t)$ by computing the rank of a $\deg_{\text{out}}(s)\times\deg_{\text{in}}(t)$ matrix. By Proposition 2.2, this takes
$\sum_{s,t\in V}\deg_{\text{out}}(s)(\deg_{\text{in}}(t))^{\omega-1}$ (15)
time asymptotically. For each pair of vertices $(s,t)$, we have
$\deg_{\text{out}}(s)(\deg_{\text{in}}(t))^{\omega-1}=(\deg_{\text{in}}(t))^{\omega-2}\cdot\deg_{\text{out}}(s)\deg_{\text{in}}(t)\leq n^{\omega-2}\cdot\deg_{\text{out}}(s)\deg_{\text{in}}(t).$
By substituting this inequality into Equation 15, and observing that the sum of in-degrees and sum of out-degrees are each equal to the number of edges $m$ in $G$, we have
$\sum_{s,t\in V}\deg_{\text{out}}(s)(\deg_{\text{in}}(t))^{\omega-1}\leq\sum_{s,t\in V}n^{\omega-2}\cdot\deg_{\text{out}}(s)\deg_{\text{in}}(t)=n^{\omega-2}m^{2}.$
Since $n-1\leq m$, the runtime of this step is also upper bounded by $\tilde{O}(m^{\omega})$.
So we can solve APC in $\tilde{O}(m^{\omega})$ time as claimed.
Algorithm 1 The algorithm solving APC from [CLL13].
1:Compute the matrix $M=(I-A)^{-1}$.
2:For each pair of vertices $(s,t)$, return
$\operatorname{rank}M[E_{\text{out}}(s),E_{\text{in}}(t)]$
as the value of $\lambda(s,t)$.
6 Bounded Connectivity
The APC algorithm from Algorithm 1 first
  1. 1.
    inverts an $m\times m$ matrix, and then
  2. 2.
    computes ranks of submatrices, whose dimensions depend on degrees of nodes in $G$.
Even reading the matrix entries in these steps takes $\Omega(m^{2})$ time. To obtain a faster algorithm for the $k$-APC problem, we modify these steps to work with much smaller matrices.
The first idea is to reduce degrees in $G$ while preserving the values of small connectivities. In Section 6.1, we present a simple transformation (from [AJ23, Section 5]) which decreases the degrees of nodes in $G$ to $k$, while preserving the $\min(k,\lambda(s,t))$ values. Following this modification, we only need to compute ranks of $k\times k$ submatrices in step 2 above.
The second idea is to simplify $\Gamma=(I-X)^{-1}$ using a variable substitution which still ensures that determinants of submatrices of $\Gamma$ can detect up to $k$ edge-disjoint paths in $G$. We present this substitution in Section 6.2. This modification turns $\Gamma$ into a rank $kn$ matrix, which lets us replace the inversion of an $m\times m$ matrix in step 1 above with the easier inversion of a $kn\times kn$ matrix instead.
These two speed-ups combined then let us solve $k$-APC in $\tilde{O}((kn)^{\omega})$ time.
6.1 Degree Reduction
Let $G$ be the input graph on $n$ nodes and $m$ edges. We modify $G$ to create a new graph.
For each vertex $v\in V$, we introduce two new nodes $v_{\text{in}}$ and $v_{\text{out}}$. Then we replace each edge $(u,v)\in E$ with an edge $(u_{\text{out}},v_{\text{in}})$. For each $v\in V$, we also include $k$ parallel edges from $v$ to $v_{\text{out}}$, and $k$ parallel edges from $v_{\text{in}}$ to $v$. Let $G_{\text{new}}$ be the new graph constructed in this way, and let $V_{\text{new}}$ and $E_{\text{new}}$ be its vertex and edge sets respectively. We refer to the nodes in $V\subseteq V_{\text{new}}$ which were originally in $G$ as the original vertices. For $s,t\in V$, we still let $\lambda(s,t)$ denote the connectivity from $s$ to $t$ in the original graph $G$. In the rest of this section, we let $E_{\text{out}}(s)$ and $E_{\text{in}}(t)$ denote the sets of edges exiting $s$ and entering $t$ in $G_{\text{new}}$.
We write $n_{\text{new}}=|V_{\text{new}}|=3n$ and $m_{\text{new}}=|E_{\text{new}}|=m+2kn$.
Lemma 6.1 (Preserving Small Connectivities).
For any $s,t\in V$, the connectivity from $s$ to $t$ in $G_{\text{new}}$ is $\min(k,\lambda(s,t))$.
Proof 6.2.
Fix $s,t\in V$. Given an $st$-path $P^{\prime}$ in $G_{\text{new}}$, we recover a unique $st$-path $P$ in $G$ by looking at the sequence of original vertices visited by $P^{\prime}$. Using this construction, any collection of $r$ edge-disjoint paths from $s$ to $t$ in $G_{\text{new}}$ recovers a collection of $r$ edge-disjoint paths from $s$ to $t$ in $G$. So the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\lambda(s,t)$.
Since $s$ has outdegree $k$ in $G_{\text{new}}$, the connectivity from $s$ to $t$ in $G_{\text{new}}$ is also at most $k$.
Thus the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at most $\min(k,\lambda(s,t))$. Set $\lambda=\min(k,\lambda(s,t))$.
By definition, there are edge-disjoint paths $P_{1},\dots,P_{\lambda}$ in $G$ from $s$ to $t$.
For each $i\in[\lambda]$, let $P^{\prime}_{i}$ be the $st$-path in $G_{\text{new}}$ which passes through the same sequence of original vertices as $P_{i}$, and includes, for each edge $(u,v)$ in $P_{i}$, the $i^{\text{th}}$ parallel edge from $u_{\text{out}}$ to $v_{\text{in}}$ (we assume there is some fixed ordering among all such parallel edges). This is possible since $\lambda\leq k$. Since the $P_{i}$ are edge-disjoint, the $P^{\prime}_{i}$ are edge-disjoint as well. So the connectivity from $s$ to $t$ in $G_{\text{new}}$ is at least $\lambda=\min(k,\lambda(s,t))$.
Thus the connectivity from $s$ to $t$ in $G_{\text{new}}$ is equal to $\min(k,\lambda(s,t))$, as claimed.
6.2 Low-Rank Edge-Adjacency
Recall the definitions from Section 6.1. We define the matrix $X$ from Section 4 with respect to the new graph $G_{\text{new}}$ (so now rows and columns of $X$ are indexed by edges in $E_{\text{new}}$).
For each pair $(e,j)\in E_{\text{new}}\times[k]$ we introduce an indeterminate $y_{ej}$.
Similarly, for each pair $(j,f)\in[k]\times E_{\text{new}}$ we introduce an indeterminate $z_{jf}$.
We define the $m_{\text{new}}\times kn_{\text{new}}$ matrix $Y$ by setting
$Y[e,(v,j)]=\begin{cases}y_{ej}&\text{if }\operatorname{head}(e)=v\\ 0&\text{otherwise.}\par\end{cases}$
Similarly, we define the $kn_{\text{new}}\times m_{\text{new}}$ matrix $Z$ by setting
$Z[(v,j),f]=\begin{cases}z_{jf}&\text{if }\operatorname{tail}(f)=v\\ 0&\text{otherwise.}\end{cases}$
These matrices are defined so that under the variable substitution
$x_{ef}=\sum_{j=1}^{k}y_{ej}z_{jf}$
the matrix $X$ simplifies to the low-rank matrix $YZ$, as depicted in Figure 2.
Previously in Corollary 4.14, we characterized the existence of edge-disjoint paths in $G$, based off whether determinants of submatrices of $(I-X)^{-1}$ were nonzero. The following result shows that a similar characterization holds when we replace $X$ with $YZ$, provided we only care about routing up to $k$ edge-disjoint paths.
$e$$f$$X[e,f]=x_{ef}$$e$$f$${\color[rgb]{0.0,0.2,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.2,0.4}(YZ)[e,f]=y_{e{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}1}}z_{{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}1}f}+y_{e{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}2}}z_{{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}2}f}+y_{e{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}3}}z_{{\color[rgb]{1.0,0.55,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{1.0,0.55,0.0}3}f}}$
Figure 2: When we substitute $x_{ef}=y_{e1}z_{1f}+\dots+y_{ek}z_{kf}$ (pictured here for $k=3$) into $X$, we get the “simpler” matrix $YZ$. While powers of $X$ enumerate walks in $G$, powers of $YZ$ intuitively enumerate walks in a modified graph where after traversing an edge $e=(u,v)$, we have $k$ different versions of $v$ we can choose to go to. The $y_{ej}$ and $z_{jf}$ variables in this enumeration only keep track of the individual edges traversed and versions of vertices we pick, instead of recording all pairs of consecutive edges traversed like the $x_{ef}$ variables. This simpler enumeration suffices to solve $k$-APC.
Lemma 6.3.
Let $S,T\subseteq E_{\text{new}}$ be subsets of edges with size $|S|=|T|=r\leq k$. Then
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]$
is a nonzero formal power series if and only if $G_{\text{new}}$ has $r$ edge-disjoint paths from $S$ to $T$.
Proof 6.4.
Suppose $G$ does not contain $r$ edge-disjoint paths from $S$ to $T$. Then by Corollary 4.14, the determinant
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]$
is identically zero as a power series. Consequently, the above expression remains zero even if we make the variable substitution
$x_{ef}=\sum_{j=1}^{k}y_{ej}z_{jf}.$ (16)
Under this substitution, the matrix $X$ simplifies to $YZ$. Thus in this case
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]$
is the zero polynomial, as claimed.
Suppose now that $G$ does not contain $r$ edge-disjoint paths from $S$ to $T$.
By Corollary 4.10, we have
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$ (17)
For each collection of walks $\mathcal{C}$, let $\tilde{\xi}(\mathcal{C})$ be the monomial resulting from substituting Equation 16 into the weight $\xi(\mathcal{C})$. Then we have
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\tilde{\xi}(\mathcal{C})\right).$ (18)
Let $\mathcal{P}=\langle P_{1},\dots,P_{r}\rangle$ be a collection of edge-disjoint paths from $S$ to $T$ in $G$.
For each $i\in[r]$, let $\mathcal{E}_{i}$ be the set of consecutive pairs of edges $(e,f)$ traversed by $P_{i}$.
Then we have
$\tilde{\xi}(\mathcal{P})=\prod_{i=1}^{r}\prod_{(e,f)\in\mathcal{E}_{i}}\left(\sum_{j=1}^{k}y_{ej}z_{jf}\right).$
If we expand the product on the right-hand side of the above equation, we see that one of the monomials produced is of the form
$\prod_{i=1}^{r}\prod_{(e,f)\in\mathcal{E}_{i}}y_{ei}z_{if}.$ (19)
Note that in this step, we are using the fact that $r\leq k$.
Because $\mathcal{P}$ is a collection of edge-disjoint paths, the variables appearing in the monomial from Equation 19 allow us to uniquely recover $\mathcal{P}$. In detail: for each index $i$, the edges $e$ for which the $y_{ei}$ variable appears recovers all edges in $P_{i}$, and because $P_{i}$ is a simple path we can recover the order of these edges as well. See Figure 3 for an example of this unique recovery property in the case of $k=2$.
Thus the monomial from Equation 19 appears with coefficient $1$. Hence
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]$
is a nonzero formal power series as claimed.
$e_{1}$$f_{1}$$e_{2}$$f_{2}$$e_{1}$$e_{2}$$f_{1}$$f_{2}$
Figure 3: Given edge-disjoint paths $P_{1}=\langle e_{1},f_{1}\rangle$ and $P_{2}=\langle e_{2},f_{2}\rangle$ in $G$, the determinant of the matrix $(I-X)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ enumerates this pair via the monomial $\xi(P_{1},P_{2})=x_{e_{1}f_{1}}\cdot x_{e_{2}f_{2}}$. The variables in this monomial provide enough information to uniquely recover $P_{1}$ and $P_{2}$. In contrast, for $k=2$, the determinant of $(I-YZ)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ assigns this pair weight $\tilde{\xi}(P_{1},P_{2})=(y_{e_{1}1}z_{1f_{1}}+y_{e_{1}2}z_{2f_{1}})(y_{e_{2}1}z_{1f_{2}}+y_{e_{2}2}z_{2f_{2}}).$ One of the terms in the expansion of the above product is $y_{e_{1}1}z_{1f_{1}}\cdot y_{e_{2}2}z_{2f_{2}}$. We can read this term as saying “the first path $P_{1}$ traverses $e_{1}$ and $f_{1}$, and the second path $P_{2}$ traverses $e_{2}$ and $f_{2}$.” So this monomial provides enough information to recover the pair of paths $\langle P_{1},P_{2}\rangle$ as well.
We also use the following lemma that simplifies computation of $(I-YZ)^{-1}$.
Lemma 6.5 (Geometric Series Identity).
We have
$(I-YZ)^{-1}=I+Y(I-ZY)^{-1}Z.$
Proof 6.6.
Since every entry of $Y$ and $Z$ have zero constant term, the same is true for the matrices $YZ$ and $ZY$. Then by the geometric series formula of Proposition 3.3, we have
$(I-YZ)^{-1}=I+(YZ)+(YZ)^{2}+\dots=I+Y\left(\sum_{\ell=0}^{\infty}(ZY)^{\ell}\right)Z.$ (20)
Applying Proposition 3.3 again, we have
$\sum_{\ell=0}^{\infty}(ZY)^{\ell}=(I-ZY)^{-1}.$
Substituting the above equation into the rightmost side of Equation 20 yields
$(I-YZ)^{-1}=I+Y(I-ZY)^{-1}Z$
as desired.
6.3 Random Evaluation
We begin by defining random evaluations of the polynomial matrices $Y$ and $Z$.
For all pairs $(e,j)\in E_{\text{new}}\times[k]$ and $(j,f)\in[k]\times E_{\text{new}}$ we introduce independent, uniform random values $b_{ej}$ and $d_{jf}$ respectively over $\mathbb{F}$. Let $B$ and $C$ be the matrices obtained from matrices $Y$ and $Z$ under the evaluations $y_{ej}=b_{ej}$ and $z_{jf}=c_{jf}$ respectively.
That is, $B$ is the $m_{\text{new}}\times kn_{\text{new}}$ matrix defined by setting
$B[e,(v,j)]=\begin{cases}b_{ej}&\text{if }\operatorname{head}(e)=v\\ 0&\text{otherwise}\par\end{cases}$
and $C$ is the $kn_{\text{new}}\times m_{\text{new}}$ matrix defined by setting
$C[(v,j),f]=\begin{cases}c_{jf}&\text{if }\operatorname{tail}(f)=v\\ 0&\text{otherwise.}\end{cases}$
We also define subsets of edges
$E_{\text{out}}=\bigcup_{s\in V}E_{\text{out}}(s)\quad\text{and}\quad E_{\text{in}}=\bigcup_{t\in V}E_{\text{in}}(t)$
and let $\tilde{B}=B[E_{\text{out}},\cdot]$ and $\tilde{C}=C[\cdot,E_{\text{in}}]$ be submatrices of $B$ and $C$ restricted to edges exiting and entering original vertices respectively.
Lemma 6.7.
Let $S,T\subseteq E_{\text{new}}$ be subsets of edges of size $|S|=|T|=r\leq k$. Then
$\det\leavevmode\nobreak\ (I-BC)^{-1}[S,T]$
is nonzero with probability at least $1-1/n^{3}$ if and only if $G_{\text{new}}$ contains $r$ edge-disjoint paths from $S$ to $T$.
Proof 6.8.
By Lemma 6.3, the formal power series
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]$ (21)
is nonzero if and only if $G_{\text{new}}$ contains $r$ edge-disjoint paths from $S$ to $T$.
If $G_{\text{new}}$ does not have $r$ edge-disjoint paths from $S$ to $T$, then Equation 21 is the zero polynomial, so its random evaluation
$\det\leavevmode\nobreak\ (I-BC)^{-1}[S,T]$
vanishes as well.
Suppose instead $G_{\text{new}}$ does contain $r$ edge-disjoint paths from $S$ to $T$. We can write
$(I-YZ)^{-1}[S,T]=\frac{(\operatorname{adj}(I-YZ))[S,T]}{\det\leavevmode\nobreak\ (I-YZ)}.$
The matrix $(I-YZ)$ has ones along its diagonal, and all its other entries have zero constant term. Hence $\det(I-YZ)$ is a polynomial with constant term $1$, so by Proposition 3.1 it has a multiplicative inverse over the ring of formal series. In particular, the above equation holds both for matrices of rational functions and formal power series.
Write $Q=\det(I-YZ)$. By the above discussion, $Q$ is a nonzero polynomial. By linearity of the determinant, we have
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\frac{\det\leavevmode\nobreak\ (\operatorname{adj}(I-YZ))[S,T]}{Q^{r}}.$ (22)
Since $Q$ is nonzero and $YZ$ is an $m_{\text{new}}\times m_{\text{new}}$ matrix where each entry has degree at most two, $Q^{r}$ has degree at most $2r\cdot m_{\text{new}}\leq 2k(m+2kn)<6n^{3}$.
Since $r$ edge-disjoint paths from $S$ to $T$ exist, Equation 18 is nonzero as a formal power series, so by Equation 22 the numerator $\det(\operatorname{adj}(I-YZ)[S,T])$ is a nonzero polynomial. Each entry of $I-YZ$ has degree at most two, so each entry of $\operatorname{adj}(I-YZ)$ has degree less than $2m_{\text{new}}$, which implies that $\det(\operatorname{adj}(I-YZ)[S,T])$ has degree at most
$2m_{\text{new}}r\leq 2(m+2kn)n\leq 6n^{3}.$
So by Corollary 2.4, the expression
$\det\leavevmode\nobreak\ (I-BC)^{-1}[S,T]$
is nonzero in this case with probability at least $1-12n^{3}/(2^{q})$. Since we picked $q$ large enough to satisfy $2^{q}\geq 12n^{6}=n^{3}\cdot(12n^{3})$, the desired result follows.
Lemma 6.9 (Small Connectivities via Rank).
With high probability, for all $s,t\in V$, we have
$\operatorname{rank}\leavevmode\nobreak\ (\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t)).$
Proof 6.10.
Fix $s,t\in V$. Let $\lambda$ be the connectivity from $s$ to $t$ in $G_{\text{new}}$.
By Lemma 6.1, $\lambda\leq k$. Thus by Lemma 6.7 and the definition of connectivity, with probability at least $1-1/n^{3}$, $\lambda$ is the largest nonnegative integer for which there exist subsets $S\subseteq E_{\text{out}}(s)$ and $T\subseteq E_{\text{in}}(t)$ of size $\lambda$ such that
$\det\leavevmode\nobreak\ (I-BC)^{-1}[S,T]$
is nonzero. In other words, $\lambda$ is equal to the rank of $(I-BC)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]$.
By applying Lemma 6.5 with the random evaluation sending $Y$ and $Z$ to $B$ and $C$ respectively, we have
$(I-BC)^{-1}=I+B(I-CB)^{-1}C.$
Since $s,t\in V$ are original vertices, we know that $E_{\text{out}}(s)\cap E_{\text{in}}(t)=\emptyset$. Thus
$(I-BC)^{-1}[E_{\text{out}}(s),E_{\text{in}}(t)]=(B(I-CB)^{-1}C)[E_{\text{out}}(s),E_{\text{in}}(t)].$
Of course we also have
$(B(I-CB)^{-1}C)[E_{\text{out}}(s),E_{\text{in}}(t)]=(\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)].$
So with probability at least $1-1/n^{3}$,
$\operatorname{rank}\leavevmode\nobreak\ (\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t))$
where we are using the fact from Lemma 6.1 that $\lambda=\min(k,\lambda(s,t))$. The claim follows by a union bound over all $n^{2}$ pairs of original vertices $(s,t)$.
6.4 The Algorithm
Algorithm 2 The algorithm solving $k$-APC from [AJ23].
1:Compute the matrix $M=\tilde{B}(I-CB)^{-1}\tilde{C}$.
2:For each pair $(s,t)$ of original vertices, return
$\operatorname{rank}M[E_{\text{out}}(s),E_{\text{in}}(t)]$
as the value for $\min(k,\lambda(s,t))$.
See 1.2
Proof 6.11.
By Lemma 6.9, Algorithm 2 correctly solves $k$-APC with high probability. It remains to bound the runtime of the algorithm.
In step 1 of Algorithm 2, we compute $\tilde{B}(I-CB)^{-1}\tilde{C}$. The matrix $CB$ has rows and columns indexed by pairs $(v,i)\in V_{\text{new}}\times[k]$. For any $(u,i),(v,j)\in V_{\text{new}}\times[k]$, we have
$CB[(u,i),(v,j)]=\sum_{e\in E_{\text{out}}(u)\cap E_{\text{in}}(v)}c_{ie}b_{ej}.$
If $v\in V$, then the summation in the right-hand side of the above equation is empty unless $u=v_{\text{in}}$, in which case the sum has exactly $k$ terms (one for each parallel edge from $v_{\text{in}}$ to $v$). So computing the entries of $CB$ corresponding to this case takes $nk^{2}\cdot k=nk^{3}$ operations.
Similarly, if $u\in V$, the sum in the above equation is empty unless $v=u_{\text{out}}$, in which case the sum consists of exactly $k$ terms. Computing the entries of $CB$ corresponding to this case takes $nk^{3}$ operations as well.
For all other cases where $u,v\in V_{\text{new}}\setminus V$, the sum consists of at most a single term. So we can compute the remaining entries of $CB$ with $((n_{\text{new}}-n)k)^{2}=4n^{2}k^{2}$ operations.
Since $k<n$ we have $nk^{3}<(nk)^{2}$, and we can compute $CB$ in $\tilde{O}((kn)^{2})$ time. Having computed $CB$, we can compute $(I-CB)^{-1}$ in $\tilde{O}((kn)^{\omega})$ time by Proposition 2.1. So step 1 of Algorithm 2 takes $\tilde{O}((kn)^{\omega})$ time overall.
Step 2 of Algorithm 2 involves computing ranks of $n^{2}$ separate $k\times k$ matrices, which by Proposition 2.2 takes $\tilde{O}(k^{\omega}n^{2})$ time.
So overall, the algorithm for $k$-APC takes $\tilde{O}((kn)^{\omega})$ time as claimed.
7 Conclusion
In this paper, we presented alternate derivations of the $\tilde{O}(m^{\omega})$ time algorithm for APC of [CLL13], and the $\tilde{O}((kn)^{\omega})$ time algorithm for $k$-APC of [AJ23]. Our approach works by testing, via random evaluation, whether certain generating functions enumerating edge-disjoint families of paths are nonzero or not.
We conclude this paper by pointing out some connections between this perspective and other classical arguments in mathematics and computer science, and highlighting the current main open problems concerning the complexity of computing connectivities.
Combinatorics
Our proof for Lemma 4.7 obtains a generating function for edge-disjoint walks by pairing up monomials corresponding to intersecting walks, and arguing their contributions cancel. This reasoning is essentially identical to the proof of the classic Lindström-Gessel-Viennot lemma from combinatorics, which is often used in mathematics to enumerate families of disjoint lattice paths. We refer the reader to [AZ18, Chapter 29] for an accessible exposition of this theorem and some of its applications. For additional examples of this technique of pairing up and cancelling extraneous terms in matrix algebra, see [Zei85].
Algebraic Algorithms
Our exposition of the APC and $k$-APC algorithms works by interpreting certain determinants combinatorially, as generating functions for families of edge-disjoint walks. Previous work has also leveraged combinatorial interpretations of the determinant (as polynomials instead of formal power series however) to construct interesting arithmetic circuits [MV97, Cur22]. In particular, the sign-reversing involution designed in [MV97, Proof of Theorem 1] is very similar to the suffix swapping argument we use in the proof of Lemma 4.7.
The general approach of solving graph problems by testing whether certain enumerating polynomials are nonzero is now a common technique in graph algorithms: we refer the reader to [EKW24] and the citations therein for previous examples of this paradigm in the literature.
Open Problems
Faster Algorithms
The main open question in this area is: can we solve APC faster? In particular, does there exist some constant $\varepsilon>0$ such that APC can be solved in general directed graphs in $O(n^{4-\varepsilon})$ time? Even obtaining such an algorithm in the special case of directed acyclic graphs would be a major breakthrough.
For constant $k$, it is conjectured that $k$-APC requires $n^{\omega-o(1)}$ time to solve, because algorithms for $k$-APC can be used to solve a problem known as Boolean Matrix Multiplication, which researchers currently do not know how to solve faster than integer matrix multiplication. Under this conjecture, the $\tilde{O}((kn)^{\omega})$ algorithm for $k$-APC is near-optimal for constant $k$. However, this does not rule out the possibility of algorithms with better dependence on $k$.
Can we obtain faster algorithms for $k$-APC when $k=n^{\delta}$ for some small constant $\delta>0$?
Better Lower Bounds
In the 4-Clique problem, we are given an undirected graph $G$ on $n$ nodes, and are tasked with determining if $G$ contains four vertices which are mutually adjacent. The best conditional lower bounds for APC come from observing that 4-Clique reduces to APC [AGI+19, Section 4]. Using the current best algorithms for rectangular matrix multiplication [WXXZ23, Table 1], the fastest known algorithm for 4-Clique takes $O(n^{3.251})$ time [EG04, Theorem 1]. Even if we conjecture this runtime time is optimal, the resulting lower bound for APC is a far cry from the $n^{4+o(1)}$ runtime we have for APC in dense graphs. Moreover, if $\omega=2$ this lower bound for APC weakens to $n^{3-o(1)}$.
Can we show better lower bounds for APC, which are supercubic even if $\omega=2$?
Faster Verification
Instead of solving APC directly, it would also be interesting to obtain better deterministic and randomized verifiers for APC. A verifier for APC is given the input to the problem, and additionally receives claims for the values of $\lambda(s,t)$ for all pairs of vertices $(s,t)$, as well as some small proof string which can be thought of as “evidence” that the claimed values are correct. The verifier reads all of these inputs, and then must determine if the claimed values are correct or not. The goal is to get a verifier which is correct (with high probability) and runs as quickly as possible.
The fastest known deterministic verifier for APC runs in $O(n^{3.251}+n^{2.5}\sqrt{m})$ time [Tra23]. It would be interesting to improve this runtime to $O(n^{3.251})$ to match the fastest known algorithm (and deterministic verifier runtime) for 4-Clique.
It would also be interesting to obtain faster randomized verifiers for APC. Currently, no randomized verifiers running faster than the deterministic verifier discussed above are known for APC. This is surprising, since the best lower bound for APC comes from the 4-Clique problem, and the 4-Clique problem admits a randomized verifier running in near-optimal $\tilde{O}(n^{2})$ time [ACJ+23, Section 4].
Does APC admit a faster randomized verifier? Alternatively, can we explain the lack of fast randomized verifiers for APC by obtaining an efficient reduction to APC from some problem which we do not believe admits fast randomized verifiers?
Vertex-Connectivity Variants
Given vertices $s$ and $t$ in graph $G$, the vertex connectivity from $s$ to $t$, denoted by $\nu(s,t)$, is the maximum number of internally vertex-disjoint paths from $s$ to $t$ in $G$. One can study the All-Pairs Vertex Connectivity (APVC) and $k$-Bounded All-Pairs Vertex Connectivity ($k$-APVC) problems as variants of APC and $k$-APVC where we are tasked with computing $\nu(s,t)$ and $\min(k,\nu(s,t))$ respectively, for all pairs $(s,t)$.
In directed graphs, APVC and $k$-APVC reduce to APC and $k$-APVC respectively, so it might be easier to design faster algorithms for the former problems instead of the latter problems. For example, it is known that $k$-APVC can be solved in $\tilde{O}(k^{2}n^{\omega})$ time [AJ23, Theorem 5], which is faster than the $\tilde{O}((kn)^{\omega})$ runtime for $k$-APC if $\omega>2$.
In undirected graphs, although APC can be solved in $\tilde{O}(n^{2})$ time, APVC is not known to admit a near-quadratic time algorithm. In fact, the best known lower bounds for APC in directed graphs also hold for APVC in undirected graphs [HLSW23].
So again, it might be easier to find faster algorithms for APVC in undirected graphs instead of tackling the general problem of APC in directed graphs. For example, APVC in undirected graphs can be solved in $\tilde{O}(m^{2})$ time [Tra23, Section 3], which is faster than the $\tilde{O}(m^{\omega})$ runtime for APC if $\omega>2$, and admits a deterministic $O(n^{3.251})$ time verifier [Tra23, Lemma 2.2], which is faster than the known $O(n^{3.521}+n^{2.5}\sqrt{m})$ time deterministic verifier for APC.
Derandomization
The algorithms we presented for APC and $k$-APC in this work are randomized. Can we solve these problems with respective $\tilde{O}(m^{\omega})$ and $\tilde{O}((kn)^{\omega})$ runtimes deterministically?